OpenStudy (anonymous):

What is the diagonal of a square with a perimeter of 48 cm?

OpenStudy (unklerhaukus):



OpenStudy (anonymous):

P=4xS 48=4x12?

OpenStudy (jiteshmeghwal9):

4s=48 s=12 area=12^2=144 \[\frac{1}{2} \times d^2= 144\]

OpenStudy (unklerhaukus):